(D+2)x-3y=1

-3x+(D+2)y=e^-t


1
Expert's answer
2022-04-07T17:21:16-0400

Multiplying first equation by 3 and solving second for 3x we’ll get

(D+2)3x-9y=3

3x = (D+2)y - e-t  

Substituting 3x from second equation into first

(D+2)[(D+2)y - e-t ] – 9y = 3

(D+2)2y + e-t – 2e-t  – 9y = 3

D2y+4Dy+4y – 9y = 3+e-t  

D2y + 4Dy – 5y = 3 +e-t  

Characteristic equation

λ2 + 4λ – 5 = 0  =>   "\lambda_{1,2}=-2\pm\sqrt9"    => λ1 = 1, λ2 = -5

So the solution of homogeneous equation is y0(t) = C1et+C2e-5t, where C1, C2 are arbitrary constants.

Partial solution may be find in the form y1(t) = A+Be-t. Substitution this into equation gives

(Be-t – 4Be-t – 5A – 5Be-t) = 3 +e-t  => -5A = 3; -8B = 1 => A = -0.6; B = -0.125

So y(t) = y0(t) + y1(t) = C1et+C2e-5t – 0.6 – 0.125e-t 

From second equation of the system

x = (1/3)[ (D+2)y - e-t] = (1/3)[ C1et – 5C2e-5t + 2C1et+2C2e-5t +0.125e-t – 1.2 – 0.25 e-t – e-t] = (1/3)[ 3C1et – 3C2e-5t – 1.2 – 1.125 e-t ] = C1et – C2e-5t – 0.4 – 0.375 e-t

Solution

x(t) = C1et – C2e-5t – 0.4 – 0.375 e-t

y(t) = C1et+C2e-5t – 0.6 – 0.125e-t